OpenStudy (anonymous):

find the slope of each line (0,0) (3,3) (-6,1) (4,8)

OpenStudy (anonymous):

i think the first one is 3/3

OpenStudy (anonymous):

m=1 m=7/10

OpenStudy (anonymous):

can u explain

OpenStudy (mwhit2101):

y2-y1 over x2-x1 is your formula for getting the slope. Plug in your values (doing your second one) \[\frac{ 8-1 }{ 4-(-6) }\] Evaluate and solve. 8-1=7 4-(-6)=10 Slope=7/10

OpenStudy (anonymous):

oh ok i get it now thank you